The price of a house decreased from $200,000 to $150,000 in one year. What is the percent decrease?
Answer by oberobic(2304)

Percentage change = (newPrice - oldPrice)/oldPrice
.
150-200/200 = -50/200
.
-50/200*100 = percent
.
-50/200*100 = -25%
.
Done.